OpenStudy (anonymous):

Please help! Solve the system of equations: 2x-5y=3 x-3y=1

OpenStudy (texaschic101):

do you prefer substitution or elimination ?

OpenStudy (anonymous):

I know substitution best.

OpenStudy (texaschic101):

x - 3y = 1 ---> x = 3y + 1 now sub 3y + 1 in for x in the other equation 2x - 5y = 3 2(3y + 1) - 5y = 3 do you know how to finish and find y ?

OpenStudy (anonymous):

No. :/

OpenStudy (texaschic101):

2(3y + 1) - 5y = 3 -- distribute through the parenthesis 6y + 2 - 5y = 3 -- subtract 2 from each side 6y - 5y = 3 - 2 -- combine like terms -y = 1 -- multiply by -1 to make y positive y = -1 now sub -1 in for y in either of the original equations x - 3y = 1 x - 3(-1) = 1 x + 3 = 1 -- subtract 3 from both sides x = 1 - 3 x = -2 check.. x - 3y = 1 -2 - 3(-1) = 1 -2 + 3 = 1 1 = 1 (correct) x = -1 and y = -2 or (-1,-2)

OpenStudy (texaschic101):

oops..x = -2 and y = -1

OpenStudy (anonymous):

Okay, so would that be the answer? Because I have these choices: (4, 1) (5, 2) (7, 2) (9, 3)

OpenStudy (texaschic101):

hmmm.....let me try something different...hold on

OpenStudy (texaschic101):

2x - 5y = 3 x - 3y = 1 --->(-2) -------------- 2x - 5y = 3 -2x + 6y = -2 (result of multiplying by -2) -------------add y = 1 now sub 1 in for y x - 3y = 1 x - 3(1) = 1 x - 3 = 1 x = 1 + 3 x = 4 I must have made a mistake on the substitution....elimination method gave me (4,1)

OpenStudy (anonymous):

Okay. Thank you so much! ^-^ Do you think you could possibly help me with 2 more questions?

OpenStudy (texaschic101):

not right now..I have to go do something...I will be on later...sorry :(

OpenStudy (anonymous):

Okay. Adios.